OpenStudy (nbarrera):

Help+Medal! Which of the following is an irrational number? A. √169 B. 0.78 C. (√12)² D. √18

OpenStudy (nbarrera):



OpenStudy (rational):

\[0.\overline{78} = \dfrac{78}{99}\] so thats NOT irrational

OpenStudy (acxbox22):

an irrational number is simply anything that cant be represented as a fraction

OpenStudy (nbarrera):



OpenStudy (nbarrera):

So 0.78 is irrational?

OpenStudy (nbarrera):

@Awolflover1

OpenStudy (nbarrera):

@Presh08

OpenStudy (nbarrera):

@acxbox22

OpenStudy (nbarrera):

do you get it

OpenStudy (acxbox22):

you can put ___ 0.78 as a fraction

OpenStudy (acxbox22):

no what is the square root of 169?

OpenStudy (acxbox22):

now*

OpenStudy (nbarrera):

answer is 13. sorry openstudy slow

OpenStudy (nbarrera):

@acxbox22

OpenStudy (acxbox22):

yes what is \[\left( \sqrt{12} \right)^2\] ?

OpenStudy (nbarrera):

12

OpenStudy (acxbox22):

correct 78/99, 13, and 12 can be represented as fractions so this leaves sqrt18 which cant be represented as a fraction therefore sqrt18 is irrational

OpenStudy (nbarrera):

Ooh. Ok I get it Thx :)